Some extesions of Eneström–Kakeya theorem

Abdul Aziz and B. A. Zargar, Srinagar, India

Keywords: Polynomials, moduli of zeros, bounds AMS Subject classification: 30C10, 30C15

Abstract: In this paper we prove several Eneström-Kakeya type theorems concerning the location of zeros of a polynomial in the complex plane. Our theorems generalise classical results by putting less restrictive conditions on the coefficients of the polynomial.